OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 @phi Can you help? I'm not good with the Quadratic Equation

OpenStudy (phi):

I think you should factor them. look at the last number: +2 1) list all pairs of numbers that give you 2 when you multiply them together in this case: 1,2 is the only pair 2) the + sign on the 2 means your "factors" will have the same sign. 3) look at the +3 in +3x the + means the largest of the factors will be + 4) from step 1, you know you have factor +2. from step 2, you know 1 is also + so your factors are +1 and +3 5) from your list of step 1 with signs as given in step 4, pick the pair that add up to +3 (this +3 comes from +3x term) in this case, there is only 1 pair, and it adds to 3. so you found the factors 6) write down (x + 1)(x+2)

OpenStudy (phi):

to solve, you need an equation. normally you solve for when y = 0 x^2 + 3x + 2 =0 if you factor (see above) you get (x+1)(x+2)=0 if x= -1 you will get 0 or if x=-2 you get 0 you would find these numbers by saying x+1=0 add -1 to both sides x+1-1= 0-1 x= -1 and x+2= 0 add -2 to both sides x+2-2= 0-2 x= -2

OpenStudy (phi):

for the 2nd equation y = x^2 + 2x + 1 can you do step 1) list all pairs of numbers that give you 1 (the last number)

OpenStudy (anonymous):

Wait, so the first one is just x=-2? and the pairs are only 1, 1.

OpenStudy (phi):

re-read the 2nd post more carefully. what does it say?

OpenStudy (anonymous):

Ohh. (x+1)(x+2)=0

OpenStudy (phi):

do you understand that has two different x values that "work" (make the equation true)

OpenStudy (anonymous):

Yes, sorry. I got distracted because I am watching a movie.

OpenStudy (phi):

if x= -1 in (x+1)(x+2)=0 you get (-1+1)*(-1+2) = 0 or 0 * (1) = 0 0 = 0

OpenStudy (anonymous):

Okay. What about the second one.

OpenStudy (phi):

or the 2nd equation y = x^2 + 2x + 1 can you do step 1) list all pairs of numbers that give you 1 (the last number) you said 1,1 now step 2) 2) the + sign on the 1 (the last number) means your "factors" will have the same sign. step 3 3) look at the +2 in +2x the + means the largest of the factors will be + what are your factors (and what sign are they) ?

OpenStudy (anonymous):

The factors are 1 and 1 right? o.o

OpenStudy (phi):

and both are +

OpenStudy (anonymous):

I'm a little confused.

OpenStudy (phi):

you got step 1. now step 2 in y = x^2 + 2x + 1 look at the last number +1 the + sign means your factors will have the same sign (either both will be + or both -) (if you had -1, it would mean the factors had opposite signs) in this case, your factors will have the same sign. Which sign ? step 3: look at the middle term in y = x^2 + 2x + 1 the + in +2x means the largest factor is +. in your case, with 1,1 pick the first 1, and say it is +1 from step 2, you know both factors have the same sign. if you have a +1 the other factor (the other 1) is also + so your factors are +1, +1

OpenStudy (phi):

when you get to a boring spot in the movie, you can read that over until it makes sense

OpenStudy (phi):

step 4 was figure out the signs: +1, and +1 step 5, from your list of pairs (in this case there is only 1 pair) pick the pair that when added together give you the middle number in x^2 + 2x + 1 you want the pair that add up to +2 (from +2x) in your case +1 + 1 =2 so those are your factors

OpenStudy (anonymous):

Okay, just so I can process this better to put it in answer form, can you help me shorten that up, so I can type it out > _ <

OpenStudy (phi):

nobody types in that much. that is for you to go over when or if you want to learn how to do it. when people "show work" they write : "factor the quadratic into ...." and write down the factored form (as if it were totally obvious to the most casual observer)

OpenStudy (anonymous):

I just needed to find a shorter way to show my work. >_<

OpenStudy (anonymous):

But my brain can't process all of the typing into a few sentences.